The number of column inches of classified advertisements appearing on Mondays in a certain daily newspaper is normally distributed with population mean of 320 and population standard deviation of 20 inches.
Referring to Scenario 6-1, for a randomly chosen Monday the probability is 0.1 that there will be less than how many column inches of classified advertisements?

Answer by stanbon(75887)

The number of column inches of classified advertisements appearing on Mondays in a certain daily newspaper is normally distributed with population mean of 320 and population standard deviation of 20 inches.
Referring to Scenario 6-1, for a randomly chosen Monday the probability is 0.1 that there will be less than how many column inches of classified advertisements?
-------
Find the z-score with a left tail of 0.1
z = invNorm(0.1) = -1.2816
-----
Find the corresponding column inches using x = z*s+u
Ans:: x = -1.2816*20+320 = 294.37 inches
-------------
Cheers,
Stan H.
--------------------